# Conditions for enhancement of chemical reactions in gas phase inside a dark cavity

### v1

@@ III.1.1 The enhancement of the rate of the reaction $H+ArCl\to[ArHCl]^{\#}\to H+Ar+Cl$
This reaction has been studied in details in Ref.20 in free space (when the distance between the two mirrors is taken to be infinitely large). In this work it has been shown that the predissociation resonances of $[ArHCl]^{\#}$ which results from the vibrations perpendicular to the reaction coordinate (when hydrogen atom oscillates between Ar and Cl) are accurately obtained from the 1D effective potential barrier that is obtained within the framework of the nuclear adiabatic approximation. Here we wish to show that in this case where all complex poles are associated with the dynamical potential barrier and are physical poles the rate of the reaction is enhanced by the dark cavity that couples the TS resonance with one of the DB resonance of $ArHCl$.
@@
Figure 1: The dynamical barrier potential (black line) and the positions (red dots) of the complex poles for ArHCl. The potential is adopted from Ref.20 and the resonance positions are calculated herein using complex scaling. Potential in Hartree and coordinate in Bohr. These resonances are actually the dynamical resonances as obtained from the full multi-dimentional calculations, notice that the static potential barrier along the reaction coordinate is negligible (see Ref.20).
@@
Figure 2: The highest five resonance complex poles of $ArHCl$, calculated by the uniform complex scaling method. All of them results from the vibrational states of $ArHCl$ when along the reaction coordinate $H$ has almost a free transnational motion with respect to the $ArCl$ center of mass. The number of nodes were counted from the plots of the square integrable complex scaled $|\Psi_{res}(X)|^{2}$. Notice that the transition state is associated the longest lifetime resonance that is localized at the top of the potential barrier. The cavity is tuned to couple the two states marked in blue. Energies in Hartree.
Below in Fig.3 we show the enhancement of the rate of this reaction when the distance between the two mirrors in the Fabrey Perot dark cavity are tuned to have a QED mode with the frequency $\omega_{cav}=0.195meV$. This QED mode corresponds to the resonance condition that couples the TS resonance and the DB resonance, which are marked in blue in Fig.2. This result encourage the experimentalists to do the suitable study in gas phase space (as for example has been carried out in Narevicius Lab in the study of cold molecular collision).
Figure 3: The rate of the reaction $H+ArCl\to H+Ar+Cl$ as a function of the coupling strength between the dark cavity and the $ArHCl$ molecule at resonance condition between the states marked in blue in Fig.2. The cavity can be tuned into this resonance condition by adjusting the distance between the two Fabrey-Perot mirrors (see Eq. 5). Detailed of the calculation are given in Section IV, in particular, Eq. 13. Rate in Hartree and coupling in atomic units.
@@ III.1.2 The enhancement of the rate of the reaction $O+D_{2}\to[ODD]^{\#}\to OD+D$
In Ref.24 Johnson and Winter studied the the effect of vibrational energy on the reaction of molecular hydrogen with atomic oxygen. They calculated the the contour map for the $O+H_{2}\xrightarrow[]{\Gamma(\epsilon_{cav}=0)}OH+H$, in which $O$ collides with $H_{2}$. See below their contour map plot as presented in Fig.1 in their paper. Notice that in the electronic structure calculations the isotopic effects are not taken into consideration and therefore we can replace hydrogen molecule by deuterium diatomic molecule. The product is $OD+D$ and the frequency of the vibration which is perpendicular to the reaction coordinate is denoted by $\Omega_{vib}(X)$. The incoming oxygen atoms collide with deuterium diatomic molecules when they are temporarily trapped for a long time around the saddle point due to the oscillations perpendicular to the reaction coordinate. This temporarily trapping results in the formation of a long lived transition state (TS) predissociation resonance. Nevertheless, there are other excited predissociation resonances which are not populated but, as we will show here, will interact with the TS resonance in a dark cavity resulting in enhancement of the reaction. The different of the masses of the isotopes will be taken into consideration in the next step where we will describe potential energy surface for this reaction using the reaction path Hamiltonian representation.
@@
The 3D and 2D potential energy surfaces as described in Fig.4 and Fig.5 were obtained respectively by using Eq.1 and Eq.2. The information for this plot comes from Johnson and Winter in 1977, see Ref.24. On this scale, one can not see the energy barrier in the X-direction. Our focus is on the weaker, ”soft” vibrations along the Y-axis, compared to the vibrations in the original reactants ($O+D_{2}$) and the products ($OD+D$). In Fig.6 one can see the static potential barrier along the reaction coordinate. As explained in Section(II) the reactants $O$ and $D_{2}$ approaching one another as they move along $X<0$. Due to the relative high frequency of the ground vibrational level of $D_{2}$ the collision energy is about equal to the height of the static potential barrier. However, as they get to be close enough to form the activated complex $[ODD]^{\#}$ the energy transfer from the excited vibrational soft mode of $[ODD]^{\#}$ to the X-component of the transnational kinetic energy will enable to bypass the potential barrier. Notice that $V_{SB}(X)$ is approximately described as an asymmetric Eckart potential barrier given by (Hartree unites) as a function pf the reaction coordinate given in Bohr unites,
@@
Figure 7: Static potential barrier as obtained from the 1D electronic structure calculations along the reaction coordinate (see Eq.3 in the text). Potential in Hartree and coordinate in Bohr.
@@
Figure 9: The adiabatic potential, obtained by adding the static potential and the dynamical barrier potential as described in the text. This adiabatic potential is exactly equal to the potential given in Eq.6.3 and in Fig.1c in Ref.25 for the $O+D_{2}\to OD+D$ reaction. Potential in Hartree and coordinate in Bohr.
By using the uniform complex scaling where the X coordinated was rotated into the complex plane by the angle $\theta=0.75rad$ [14] the resonances for the adiabatic Hamiltonian for $OD_{2}$ were calculated and were counted by the number of their nodes. The results for the transition state (TS) and dynamical barrier (DB) resonances are given in Fig.10. By comparing the static and the adiabatic potentials as are presented in Fig.7 and Fig.9 it is clear that the the complex poles presented in Fig.10 are all associated with the dynamical barrier potential and not with the static potential barrier. Notice that the transition state is associated the longest lifetime resonance that is localized at the top of the adiabatic potential barrier ($\approx 0.0015$ hartree). The other resonances (denoted as dynamical potential-barrier resonances) are associated with the excited vibrations which are perpendicular to the reaction coordinate. The cavity is tuned to couple the two states marked in blue in Fig.10.
Figure 10: The highest six complex poles of $[ODD]^{\#}$, calculated by the uniform complex scaling method. Energies in Hartree.
The ability to enhance the rate of the reaction $O+D_{2}\to OD+D$ is due to the fact that there is a DB resonance (mark by a blue color) which its position is lower than the position of the TS resonance (mark also by a blue color). Therefore through emission of one photon the TS resonance will have the same energy as of the five nodes excited DB resonance. As we will show below in the next section the rate of this reaction is enhanced when the distance between the two mirrors is tuned to have a QED state with the energy $\hbar\omega_{cav}$ which is equal to the energy gap between the TS and DB resonances. That is, $\hbar\omega_{cav}=Re[Eres(TS)-Eres(DB)]=0.0001547Hartree$. The condition that $Re[Eres(TS)-Eres(DB)]>0$ implies that due to the emission of one photon the two resonances becomes almost degenerate with the energy of the DB resonance (this resembles the situation in standard stimulated emission). In the next section we will describe how the two molecular states are mixed with the QED waves to form upper and lower polartions. This condition should be satisfied in any other reaction that its rate can be enhanced by a dark cavity.
@@ IV The enhancing of the reaction rate inside a dark cavity : theory and example
In scattering calculations, as in the study of chemical reactions, it is preferable to use the acceleration guage representation since the dipole transition matrix elements vanish at the asymptotes. Within the acceleration gauge representation[26] the cavity induces a dipole interaction between the poles given by $0.5\alpha_{cav}\langle TS|\partial V_{ad}/\partial X|DB\rangle$, where $\alpha_{cav}=\epsilon_{cav}/\omega_{cav}^{2}$ (see Ref.26) are the cavity parameters. Therefore, in the acceleration guage representation
@@
$[-\frac{1}{2}(\partial_{X}^{2}+\partial_{Y}^{2})+V(X,Y)]|0_{photon}\rangle\langle 0_{photon}|+$
$[-\frac{1}{2}(\partial_{X}^{2}+\partial_{Y}^{2})+V(X,Y)+\hbar\omega_{cav}]|1_{photon}\rangle\langle 1_{photon}|+$
@@
where $\alpha_{0}=\epsilon_{cav}/\omega_{cav}^{2}$ and $X,Y$ are mass weighted coordinates. Here we are coming to the association of $\alpha_{0}$ with the length of the distance between the two mirrors denoted by $L$.
@@
where A is the area of the mirrors and $LA$ is the volume of the cavity. It is clear that when $L\to\infty$ the strength of the coupling between the molecules and the cavity is reduced to zero. When $L=\lambda/2$ where $\omega_{cav}=4\pi c/L$ we get that $L=4\pi c/\omega_{cav}$. Only when $\hbar\omega_{cav}$ is equal to the difference between the TS energy and the energy of the selected DB resonance the rate of the reaction will be enhanced by the cavity as we will show here.
@@
Notice that, $E_{DB}+\hbar\omega_{cav}=E_{TS}$. In dark cavity where the energy of TS resonance is larger than the energy of the DB resonance it implies that the TS emits a photon with the energy $\hbar\omega_{cav}$. The coupling between them is induced by the cavity QED modes to create a superposition of the two polaritons where the decay rate of the TS resonance is increased while the rate of the DB resonance is reduced. This mechanism is described in more details below.
@@ IV.1 Theory - the cavity effect on the rate of a reaction
The interaction of the activated complex in its transition state configuration with the planar cavity QED waves is described by a dressed Hamiltonian. We use as a basis set to construct the dressed Hamiltonian a mixture of molecular and QED states (so called polaritons). The minimal basis set for the AC outside of the cavity is constructed from the Siegert complex solutions described in the previous paragraph, see for example in Ref.27. We use for the study of the reactions inside the cavity two polariton basis functions. One polariton basis function which describes the initial state (TS outside of the cavity) is $|TS\rangle|0_{photon}\rangle$. The second basis function is $|TS\rangle|1_{photon}\rangle$ which is coupled to the initial state by the cavity. The $|TS\rangle$ is also a complex Siegert metastable solution of the $H+CH_{4}$, which is associated with the complex eigenvalue $E_{TS}-i\Gamma_{TS}/2$. Therefore, on resonance condition where
@@
(i.e., due to the coupling between the molecules modes and QED wave a photon is emitted) and
@@
Often the Hamiltonian of the polaritons is written within the length guage representation. However, in scattering calculations, as in the study of chemical reactions, it is preferable to use the acceleration guage representation since the dipole transition matrix elements vanish at the asymptotes. Therefore, in the acceleration guage representation
$\hat{H}_{polariton}=$ (8)
$[-\frac{1}{2}(\partial_{X}^{2}+\partial_{Y}^{2})+V(X,Y)]|0_{photon}\rangle\langle 0_{photon}|+$
$[-\frac{1}{2}(\partial_{X}^{2}+\partial_{Y}^{2})+V(X,Y)+\hbar\omega_{cav}]|1_{photon}\rangle\langle 1_{photon}|+$
where $\alpha_{0}=\epsilon_{cav}/\omega_{cav}^{2}$ and $X,Y$ are mass weighted coordinates. A $2\times 2$ complex symmetric matrix representation of Eq.8 in the field free resonance basis set is given by Eq. 6. Notice that, $E_{DB}+\hbar\omega_{cav}=E_{TS}$, in addition, in scattering calculations it is preferable to use the acceleration gauge representation (see for example Ref.28).
@@
Emission of a photon with the energy of $\hbar\omega_{cav}$ enables the coupling between the TS and the the DB resonances. Upon the situation that the discriminant is vanished the rate of the equation is reduced since $\Gamma_{TS}<<\Gamma_{DB}$. However as $\epsilon_{cav}$ is increased (and therefore also $\alpha_{cav}$) then the discriminant is a positive number when the dipole transition is real, and consequently the upper polariton decays approximately as at the EP condition. This result, where the decay rate is not increased when the intensity of the coupling of the activated complex with the QED modes is increased, reminds of how in 1933 Bethe explained the unusual long lifetime of hydrogen in $2s$ state in strong dc field (see Eq. 25 in Ref. 29). In our case the dipole transitions is complex and therefore
$\Gamma_{min}^{polaritation}=\frac{\Gamma_{TS}+\Gamma_{DB}}{4}-2\sqrt{\alpha^{2}_{cav}Im[d_{{TS,DB}}^{2}]}$
@@
$\Gamma_{min,\pm}^{polaritation}=\frac{\Gamma_{TS}+\Gamma_{DB}}{4}\pm 2\sqrt{\frac{\Re[d^{2}_{{TS,DB}}]Im[d_{{TS,DB}}^{2}]}{(\Gamma_{TS}-\Gamma_{DB})^{2}}}$ (10)
The calculations of the effect on the cavity on on the reaction rate is carried out under the approximation that the cavity frequency $\omega_{cav}$ is tuned to couple the TS complex pole (i.e., TS- resonance) and one of the dynamical-barrier complex poles, so called DB-resonance. See Fig.10. This approximation holds when the strength coupling parameter between the molecular system and cavity is weak. When the complex poles are used as a basis set the polaritons (mixture of molecular states and QED states; $|E_{res}^{TS}-\frac{i}{2}\Gamma_{res}^{TS}\rangle|0_{photon}\rangle$, $|E_{res}^{DB}-\frac{i}{2}\Gamma_{res}^{DB}\rangle|1_{photon}\rangle$) the following time-independent Schrödeinger equation is obtained,

### v3

@@ III.1.1 The enhancement of the rate of the reaction $H+ArCl\to[ArHCl]^{\#}\to H+Ar+Cl$
This reaction has been studied in details in Ref.19 in free space (when the distance between the two mirrors is taken to be infinitely large). In this work it has been shown that the predissociation resonances of $[ArHCl]^{\#}$ which results from the vibrations perpendicular to the reaction coordinate (when hydrogen atom oscillates between Ar and Cl) are accurately obtained from the 1D effective potential barrier that is obtained within the framework of the nuclear adiabatic approximation. Here we wish to show that in this case where all complex poles are associated with the dynamical potential barrier and are physical poles the rate of the reaction is enhanced by the dark cavity that couples the TS resonance with one of the DB resonance of $ArHCl$. The physical poles are associated with the Feshbach resonances of the 2D potential energy surface, and their associated energies (the real part of the complex energy of the pole of the scattering matrix) are larger than the threshold energy of the 1D effective potential. All other poles are non-physical and can only be used as part of the discrete basis set constructed from all complex poles of the scattering matrix. A simple distinction between the physical and non-physical complex poles can be made by comparing the complex poles associated with shape-type resonances obtained from the 1D potential as a function of the reaction coordinate against the complex poles of the adiabatic effective 1D potential energy curve. The complex poles of the adiabatic 1D potential energy curve approximate the 2D Feshbach resonances mentioned above. The poles associated with the 1D shape-type resonances are false resonances and they are very different from the Feshbach resonances obtained from the 2D reaction path Hamiltonian, which are considered the physical complex poles.
@@
Figure 1: The dynamical potential barrier (black line) and the positions (red dots) of the complex poles for ArHCl. The potential is adopted from Ref.19 and the resonance positions are calculated herein using complex scaling. Potential in Hartree and coordinate also in atomic units. These resonances are actually the dynamical resonances as obtained from the full multi-dimensional calculations, notice that the static potential barrier along the reaction coordinate is negligible (see Ref.19).
@@
Figure 2: The highest five resonance complex poles of $ArHCl$, calculated by the uniform complex scaling method. All of them results from the vibrational states of $ArHCl$ when along the reaction coordinate $H$ has almost a free translational motion with respect to the $ArCl$ center of mass. The number of nodes were counted from the plots of the square integrable complex scaled $|\Psi_{res}(X)|^{2}$. Notice that the transition state is associated the longest lifetime resonance that is localized at the top of the potential barrier. The cavity is tuned to couple the two states marked in blue. Energies in Hartree.
Below in Fig.3 we show the enhancement of the rate of this reaction when the distance between the two mirrors in the Fabry Perot dark cavity are tuned to have a quantized field mode with the frequency $\omega_{cav}=0.195meV$. This quantized field mode corresponds to the resonance condition that couples the TS resonance and the DB resonance, which are marked in blue in Fig.2.
This result encourages experimentalists to study such phenomena in the context of either cold or room temperature molecular collisions in gas phase, as for example in Narevicius Lab[25]. Generally speaking, the relevant temperatures should be high enough as to allow overcoming classically the static potential barrier. In the particular case discussed here (the collision of deuterium diatomic molecules with oxygen) the experiment has to be done at room temperature. On the other hand, an experiment involving $ArHCl$ should be done preferably at cold temperatures, because the static potential barrier along the reaction coordinate is negligible. In the experiments of Ref.25, controlling the temperature implies to increase the angle between the angle of the two colliding atomic/molecular beams (this is equivalent to increase the temperature of the collision).
Figure 3: The rate of the reaction $H+ArCl\to H+Ar+Cl$ as a function of the coupling strength between the dark cavity and the $ArHCl$ molecule at resonance condition between the states marked in blue in Fig.2. The cavity can be tuned into this resonance condition by adjusting the distance between the two Fabry-Perot mirrors (see Eq. 6). Detailed of the calculation are given in Section IV, in particular, Eq. 13. Rate in Hartree and coupling in atomic units. Notice that $E_{TS}-E_{DB}(3nodes)\approx\hbar\omega_{cav}(1_{photon}+\frac{1}{2})$. Here we used such a distance between the two mirrors that generates one mode resonant with the molecular transition (i.e., $n_{photon}=1$) as mentioned in Ref.13. The enhancement is saturated quickly, since in this case the probabilities of the upper and lower polaritons to populate the transition state become about equal as $\epsilon_{cav}$ is increased. Roughly speaking, the rate of reaction is about equal to the averaged value of the TS decay rate outside of the cavity and the decay rate of the broad three mode resonance (shown in Fig.2). This point will be explained later by Eq.12.
@@ III.1.2 The enhancement of the rate of the reaction $O+D_{2}\to[ODD]^{\#}\to OD+D$
In Ref.26 Johnson and Winter studied the the effect of vibrational energy on the reaction of molecular hydrogen with atomic oxygen. They calculated the the contour map for the $O+H_{2}\xrightarrow[]{\Gamma(\epsilon_{cav}=0)}OH+H$, in which $O$ collides with $H_{2}$. See below their contour map plot as presented in Fig.1 in their paper. Notice that in the electronic structure calculations the isotopic effects are not taken into consideration and therefore we can replace hydrogen molecule by deuterium diatomic molecule. The product is $OD+D$ and the frequency of the vibration which is perpendicular to the reaction coordinate is denoted by $\Omega_{vib}(X)$. The incoming oxygen atoms collide with deuterium diatomic molecules when they are temporarily trapped for a long time around the saddle point due to the oscillations perpendicular to the reaction coordinate. This temporarily trapping results in the formation of a long lived transition state (TS) predissociation resonance. Nevertheless, there are other excited predissociation resonances which are not populated but, as we will show here, will interact with the TS resonance in a dark cavity resulting in enhancement of the reaction. The different of the masses of the isotopes will be taken into consideration in the next step where we will describe potential energy surface for this reaction using the reaction path Hamiltonian representation.
@@
The 3D and 2D potential energy surfaces as described in Fig.4 and Fig.5 were obtained respectively by using Eq.1 and Eq.2. The information for this plot comes from Johnson and Winter in 1977, see Ref.26. On this scale, one can not see the energy barrier in the X-direction. Our focus is on the weaker, ”soft” vibrations along the Y-axis, compared to the vibrations in the original reactants ($O+D_{2}$) and the products ($OD+D$). In Fig.6 one can see the static potential barrier along the reaction coordinate. As explained in Section(II) the reactants $O$ and $D_{2}$ approaching one another as they move along $X<0$. Due to the relative high frequency of the ground vibrational level of $D_{2}$ the collision energy is about equal to the height of the static potential barrier. However, as they get to be close enough to form the activated complex $[ODD]^{\#}$ the energy transfer from the excited vibrational soft mode of $[ODD]^{\#}$ to the X-component of the translational kinetic energy will enable to bypass the potential barrier. Notice that $V_{SB}(X)$ is approximately described as an asymmetric Eckart potential barrier given by (Hartree unites) as a function pf the reaction coordinate given in Bohr unites,
@@
Figure 7: Static potential barrier as obtained from the 1D electronic structure calculations along the reaction coordinate (see Eq.3 in the text). Potential in Hartree and coordinate in Bohr. Notice that the energy of the reactants, $O+D_{2}$, is taken as zero.
@@
Figure 9: The adiabatic potential, obtained by adding the static potential and the dynamical barrier potential as described in the text. This adiabatic potential is exactly equal to the potential given in Eq.6.3 and in Fig.1c in Ref.27 for the $O+D_{2}\to OD+D$ reaction. Potential in Hartree and coordinate in atomic unites.
By using uniform complex scaling, where the X coordinate was rotated into the complex plane by an angle of $\theta=0.75rad$ [16], the resonances for the adiabatic Hamiltonian for $OD_{2}$ were calculated and counted by the number of their nodes. The results for the transition state (TS) and dynamical barrier (DB) resonances are given in Fig. 10. By comparing the static and adiabatic potentials presented in Fig. 7 and Fig. 9, it is clear that the complex poles in Fig. 10 are all associated with the dynamical barrier potential and not with the static potential barrier. Notice that the transition state is associated with the longest lifetime resonance, which is localized at the top of the adiabatic potential barrier (approximately 0.0015 hartree). The other resonances (denoted as dynamical potential-barrier resonances) are associated with the excited vibrations that are perpendicular to the reaction coordinate. The cavity is tuned to couple the two states marked in blue in Fig. 10.
Figure 10: The highest six complex poles of $[ODD]^{\#}$, calculated by the uniform complex scaling method. Energies in Hartree. The two resonances that will be coupled by the cavity are marked in blue.
The ability to enhance the rate of the reaction $O+D_{2}\to OD+D$ is due to the fact that there is a DB resonance (mark by a blue color) which its position is lower than the position of the TS resonance (mark also by a blue color). Therefore through emission of one photon the TS resonance will have the same energy as of the five nodes excited DB resonance. As we will show below in the next section the rate of this reaction is enhanced when the distance between the two mirrors is tuned to have a quantized field mode state with the energy $\hbar\omega_{cav}$ which is equal to the energy gap between the TS and DB resonances. That is, $\hbar\omega_{cav}=Re[Eres(TS)-Eres(DB)]=0.0001547Hartree$. The condition that $Re[Eres(TS)-Eres(DB)]>0$ implies that due to the emission of one photon the two resonances becomes almost degenerate with the energy of the DB resonance (this resembles the situation in standard stimulated emission). In the next section we will describe how the two molecular states are mixed with the quantized radiation field modes to form upper and lower polaritons. This condition should be satisfied in any other reaction that its rate can be enhanced by a dark cavity.
@@ IV The enhancing of the reaction rate inside a dark cavity : theory and example
In scattering calculations, as in the study of chemical reactions, it is preferable to use the acceleration gauge representation since the dipole transition matrix elements vanish at the asymptotes. Within the acceleration gauge representation[28] the cavity induces a dipole interaction between the poles given by $0.5\alpha_{cav}\langle TS|\partial V_{ad}/\partial X|DB\rangle$, where $\alpha_{cav}=\epsilon_{cav}/\omega_{cav}^{2}$ (see Ref.28) are the cavity parameters. Therefore, in the acceleration gauge representation
@@
$[\hat{T}_{X}+\hat{T}_{Y}+V(X,Y)]|0_{photon}\rangle\langle 0_{photon}|+$
$[\hat{T}_{X}+\hat{T}_{Y})+V(X,Y)+\hbar\omega_{cav}]|1_{photon}\rangle\langle 1_{photon}|+$
@@
where $V(X,Y)$ is the potential as obtained from the RPH, $\alpha_{0}=\epsilon_{cav}/\omega_{cav}^{2}$, $X,Y$ are the mass weighted coordinates, and $T_{X}=-(1/2)\partial_{XX}$, $T_{Y}=-(1/2)\partial_{YY}=-(1/2)\sum_{j=1}^{3N-7}\partial_{Y_{j}Y_{j}}$. For the sake of clarity, let us point out that the Hamiltonian of Eq. 4 corresponds to the standard reaction path Hamiltonian as derived in Ref.29. Here $X$ stands for the reaction coordinate. The used coordinate system $(X,Y)$ is curvilinear, and this implies a nontrivial form of the kinetic energy operator. In Eq. 4 we have however used a simplified approximated version of the kinetic energy operator, which neglects the Coriolis effects. Such an approximation is justified provided that the reaction takes place in close vicinity of the saddle point.
Within the adiabatic approximation the polariton Hamiltonian is given by,
$\hat{H}^{ad}_{polariton}=$ (5)
$[-\frac{1}{2}\partial_{XX}+V_{r.c.}(X)+\hbar\Omega(X)/2]|0_{photon}\rangle\langle 0_{photon}|+$
$[\hat{T}_{X}+V_{r.c.}(X)+\hbar\Omega(X)/2+\hbar\omega_{cav}]|1_{photon}\rangle\langle 1_{photon}|+$
An analogous kind of adiabatic approximation has been used before in the context of the reaction path Hamiltonians e.g. in Ref.30 (see Eq. 10 therein). The TS and DB resonances are obtained by solving Eq.5 by using the complex scaling approach[16].
Here we are coming to the association of $\alpha_{0}$ with the length of the distance between the two mirrors denoted by $L$.
@@
where A is the area of the mirrors and $LA$ is the volume of the cavity. It is clear that when $L\to\infty$ the strength of the coupling between the molecules and the cavity is reduced to zero. When $L=\lambda/2$ where $\omega_{cav}=4\pi c/L$ we get that $L=4\pi c/\omega_{cav}$. Only when $\hbar\omega_{cav}$ is equal to the difference between the TS energy and the energy of the selected DB resonance the rate of the reaction will be enhanced by the cavity as we will show here. Notice that, although we discuss here the interaction of a single molecule with the cavity, our presented theory holds also when there are $N$ identical molecules inside the cavity that do not interact with each other (but all of them interact with the same dark cavity mode). In such a case, $\epsilon_{cav}$ for a single molecule is multiplied by $\sqrt{N}$ (see Ref.31).
@@
Notice that, $E_{DB}+\hbar\omega_{cav}=E_{TS}$. In dark cavity where the energy of TS resonance is larger than the energy of the DB resonance it implies that the TS emits a photon with the energy $\hbar\omega_{cav}$. The coupling between them is induced by the cavity quantized field modes to create a superposition of the two polaritons where the decay rate of the TS resonance is increased while the rate of the DB resonance is reduced. This mechanism is described in more details below.
@@ IV.1 Theory - the cavity effect on the rate of a reaction
The interaction of the activated complex in its transition state configuration with quantized radiation field modes of a planar cavity is described by a dressed Hamiltonian. We use as a basis set to construct the dressed Hamiltonian a mixture of molecular and quantized field mode states (so called polaritons). The minimal basis set for the AC outside of the cavity is constructed from the Siegert complex solutions described in the previous paragraph, see for example in Ref.32. We use for the study of the reactions inside the cavity two polariton basis functions. One polariton basis function which describes the initial state (TS outside of the cavity) is $|TS\rangle|0_{photon}\rangle$. The second basis function is $|TS\rangle|1_{photon}\rangle$ which is coupled to the initial state by the cavity. The $|TS\rangle$ state is also metastable (a Siegert type eigensolution) if the activated complex is associated with the complex eigenvalue $E_{TS}-i\Gamma_{TS}/2$. Therefore, on resonance condition where
@@
(i.e., due to the coupling between the molecules modes and quantized field mode a photon is emitted) and
@@
Often the Hamiltonian of the polaritons is written within the length gauge representation. However, in scattering calculations, as in the study of chemical reactions, it is preferable to use the acceleration gauge representation since the dipole transition matrix elements vanish at the asymptotes.See Eq.4 for the Hamiltonian for polaritons in the acceleration gauge representation. A $2\times 2$ complex symmetric matrix representation of Eq.4 in the field free resonance basis set is given by Eq. 7. Notice that, $E_{DB}+\hbar\omega_{cav}=E_{TS}$, in addition, in scattering calculations it is preferable to use the acceleration gauge representation (see for example Ref.33).
@@
Emission of a photon with the energy of $\hbar\omega_{cav}$ enables the coupling between the TS and the the DB resonances. Upon the situation that the discriminant is vanished the rate of the equation is reduced since $\Gamma_{TS}<<\Gamma_{DB}$. However as $\epsilon_{cav}$ is increased (and therefore also $\alpha_{cav}$) then the discriminant is a positive number when the dipole transition is real, and consequently the upper polariton decays approximately as at the exceptional point (EP) condition. The EP corresponds to a non Hermitian degeneracy (see chapter 9 in Ref.16). Such a non-Hermitian degeneracy occurs when the discriminant in Eq.9 vanishes, and the decay rate of the polariton is about equal to the average value of the reaction rate in the transition state and of the excited state which decays faster. This result, where the decay rate is not increased when the intensity of the coupling of the activated complex with the quantized field modes is increased, reminds of how in 1933 Bethe explained the unusual long lifetime of hydrogen in $2s$ state in strong dc field (see Eq. 25 in Ref. 34). In our case the dipole transitions is complex and therefore
$\Gamma_{min}^{polariton}=\frac{\Gamma_{TS}+\Gamma_{DB}}{4}-2\sqrt{\alpha^{2}_{cav}Im[d_{{TS,DB}}^{2}]}$
@@
$\Gamma_{min,\pm}^{polariton}=\frac{\Gamma_{TS}+\Gamma_{DB}}{4}\pm 2\sqrt{\frac{\Re[d^{2}_{{TS,DB}}]Im[d_{{TS,DB}}^{2}]}{(\Gamma_{TS}-\Gamma_{DB})^{2}}}$ (10)
The calculations of the effect on the cavity on on the reaction rate is carried out under the approximation that the cavity frequency $\omega_{cav}$ is tuned to couple the TS complex pole (i.e., TS- resonance) and one of the dynamical-barrier complex poles, so called DB-resonance. See Fig.10. This approximation holds when the strength coupling parameter between the molecular system and cavity is weak. When the complex poles are used as a basis set the polaritons (mixture of molecular states and quantized field mode states; $|E_{res}^{TS}-\frac{i}{2}\Gamma_{res}^{TS}\rangle|0_{photon}\rangle$, $|E_{res}^{DB}-\frac{i}{2}\Gamma_{res}^{DB}\rangle|1_{photon}\rangle$) the following time-independent Schrödeinger equation is obtained,